L: R3 to R3 defined by L([a1 a2 a3]) = [a1 a2^2+a3^2 a3^2].
Prove that this is a linear transformation or not.
Answer #1

L(A + B) = L(a1 + b1, a2 + b2, a3 + b3) = [a1 + b1, (a2 + b2)2 + (a3 + b3)2, (a3 + b3)2]

= [a1 + b1, a22 + b22 + 2a2b2 + a32 + b32 + 2a3b3, a32 + b32 + 2a3b3]

= [a1, a22 + a32, a32] + [b1, b22 + b32, b32] + [0, 2a2b2, 2a3b3]

= L(A) + L(B) + [0, 2a2b2, 2a3b3] ≠ L(A) + L(B)

therefore L is not a linear transformation since:

L(A+B) ≠ L(A) + L(B)